state the slope of the graph of f.
f(x)=9-x

Answer by lwsshak3(11628)

state the slope of the graph of f.
f(x)=9-x
This is an equation of a straight line: y=mx+b, m=slope, b=y-intercept
rewrite given equation:
y=-x+9
slope=-1
y-intercept=9